OpenStudy (anonymous):

need help the graph of square root of x has symmetry , but the graph of -square root of x does not A.true B.false

OpenStudy (anonymous):

mathmate can u help me

OpenStudy (mathmate):

Can you find me a function which has symmetry and the reflection (-f(x)) doe not?

OpenStudy (anonymous):

wait hold on i forgot to post the other question hang on

OpenStudy (anonymous):

the graph of y=1/x satisfies which of the following A.it intersects the x-axis B.it is symmetric to the origin C.it intersects the y-axis D.the end behavior of the function is that it approaches zero Select all correct answers

OpenStudy (mathmate):

@Twilliams250 Are you back?

OpenStudy (anonymous):

yea

OpenStudy (mathmate):

Can you find me a function which has symmetry and the reflection (-f(x)) doe not?

OpenStudy (anonymous):

i need the question i just posted i got the first one already

OpenStudy (mathmate):

What did you get for the response?

OpenStudy (anonymous):

i googled it and found out it was true

OpenStudy (mathmate):

How so?

OpenStudy (mathmate):

can you show me the page?

OpenStudy (anonymous):

can we work on the problem first i need help with it

OpenStudy (anonymous):

the graph of y=1/x satisfies which of the following A.it intersects the x-axis B.it is symmetric to the origin C.it intersects the y-axis D.the end behavior of the function is that it approaches zero Select all correct answers

OpenStudy (mathmate):

Hint: 

OpenStudy (anonymous):

can u help me solve it

OpenStudy (anonymous):

@aum can u help me

OpenStudy (aum):

y=1/x is symmetric about the origin as shown by mathmate's diagram above. The first one is false. The square root of x has no symmetry.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@aum

OpenStudy (aum):

@Twilliams250

OpenStudy (anonymous):

is b,c,d correct

OpenStudy (aum):

y=1/x is symmetric about the origin. That is the only correct answer. The others are false.

OpenStudy (aum):

B is the correct one. A,C, D are false.

OpenStudy (anonymous):

can u help me with another one

OpenStudy (aum):

Post a separate question. Since you posted this same question again you can probably edit that.

OpenStudy (aum):

I will have to mention that the "end behavior" choice is a bit ambiguous if you look at the graph in the first quadrant, one end approaches zero while the other end approaches +infinity. In the third quadrant, one end approaches zero while the other end approaches -infinity. So if the "end behavior" is defined as x approaching plus or minus infinity, then D is correct. But if "end behavior" includes points of discontinuity, then D is false.

OpenStudy (anonymous):

the reciprocal function f(x)=1/x satifies wich of the following A.it exibits direct variation B.it exibits inverse variation C.its range is all real numbers D.its domain is all real numbers except 0

OpenStudy (aum):

Direct variation is f(x) = k * x where k is a constant. Inverse variation is f(x) = k / x. So which one do you think is correct between A and B?

OpenStudy (anonymous):

a

OpenStudy (aum):

Not A. It is inverse variation with k = 1. f(x) = 1 / x

OpenStudy (anonymous):

oh ok